OpenStudy (tomektomeknyc):

if arc tan x + 2 Arc tan 1 = 3 pi/4, tan x = ? a) 1 b) -1 c) 0 d) pi/4

OpenStudy (raden):

Arc tan 1 = pi/4, right ?

OpenStudy (tomektomeknyc):

its 3 pi/4 in a question it mi ght be a typo though

OpenStudy (raden):

i mean, the value of Arc tan 1 = pi/4. because tan(pi/4) = 1, agree ? so, arc tan x + 2 Arc tan 1 = 3 pi/4 arc tan x + 2 pi/4 = 3 pi/4 arc tan x = 3 pi/4 - 2pi/4 arc tan x = pi/4 now put tan to both sides, get tan (arc tan x) = tan(pi/4) x = 1 are you sure the question is find tan(x), not only x ?

OpenStudy (tomektomeknyc):

yes, only tan x Thank you very much

OpenStudy (raden):

wlcm